OpenStudy (anonymous):

There are 33 students in the club. There are five more boys than girls in the club. Write and solve a system of equations to find the number of boys and girls in the club. Solve by graphing.

OpenStudy (pfenn1):

if x=number of boys and y= number of girls and the total number of students is 33 then \[x+y=33\] If there are 5 more boys than girls, how would you write that in an equation?

OpenStudy (anonymous):

(x+5) + y =33

OpenStudy (pfenn1):

Very close. 5 more boys than girls means that the number of boys equals the number of girls plus 5. so\[x=y+5\]So our two equations are: \[x=y+5\]\[x+y=33\] Plug the expression for x into the second equation and solve for y.

OpenStudy (anonymous):

ok - how would I put tht on a graph?

OpenStudy (pfenn1):

You have to graph it?

OpenStudy (anonymous):

I think so, there is a blank graph next to the question.

OpenStudy (pfenn1):

what did you get for x and y?

OpenStudy (anonymous):

x=19 and y=14

OpenStudy (pfenn1):

So the only thing you can graph is a point where x=19 and y=14, isn't it? Or do they want you to graph each of the equations and that point is where the two lines would cross?

OpenStudy (anonymous):

I think they want a graph with each of the equations - two lines